OpenStudy (anonymous):

Subtract... 9/16-)3/16)

OpenStudy (anonymous):

(9-3)/16=5/16

jhonyy9 (jhonyy9):

this is not correct 9/16 - 3/16 = (9-3)/16 = 6/16 = 3/8